Determine whether the system of linear equations has one and only one solution, infinitely many solutions, or no solution. Find all solutions whenever they exist.
x + 2y = 4
2x - y = 3

Answer by mananth(16946)

x+2y=4 .............1
2x-y=3 .............2

multiply (1)by 1
Multiply (2) by 2
1x+2y=4
4x-2y=6

Add the two equations

5x=10

/5

x=2

plug value of x in (1)

x+2y=4
2+2y=4
2y=4-2
2y=2
y=1